OpenStudy (anonymous):

expand function f(z)= (z^2)-1 in Taylor series at z=1

OpenStudy (amistre64):

f(z) = z^2 -1 f'(z) = 2z f''(z) = 2 f'''(z) = 0

OpenStudy (amistre64):

is it to the (z-1)^n or (z+1)^n ??

OpenStudy (anonymous):

Thanks. It is (z^2)-1 . as you did the function is not differentiable after f''' (z)

OpenStudy (amistre64):

it is ...but its pointless to add up alot of zeros ;)

OpenStudy (anonymous):

Thanks . still think about it :)

OpenStudy (amistre64):

0 2 2 --- + --- (x+1) + --- (x+1)^2 ...this it? 0! 1! 2!

OpenStudy (amistre64):

...(x-1)s right?

OpenStudy (amistre64):

2x -2 +x^2 -2x +1 x^2 -1...if im seeing it right lol

OpenStudy (anonymous):

Is it \[z^2-1?\] OR \[{(z^2)}^{-1}?\]

OpenStudy (anonymous):

z ^2 −1

OpenStudy (anonymous):

at z=1

OpenStudy (amistre64):

....i take it that formula subs z for x then lol

OpenStudy (amistre64):

so after we expand it; we apply z=1 and get 0 right?

OpenStudy (amistre64):

i just started readig taylors the other day :)

OpenStudy (anonymous):

Yeah, I think what you did is right.

OpenStudy (anonymous):

It just should be (x-1) rather than (x+1). Right?

OpenStudy (amistre64):

yes...(x-1) :) the general is (x-a) so i assumed a=1

OpenStudy (anonymous):

Exactly. It's z, not x anyways :)

OpenStudy (anonymous):

Thanks guys :)